Expedia and Priceline compete as online travel agencies. Historically, Expedia has focused more on flights, whereas Priceline has focused on hotel bookings. The following amounts were reported by the two companies in 2015 Net Income Total ssets Liabilities Revenues Total Total (in millions) $723 $15,504 Expedia Priceline $10,574 8,625 $6,672 9,224 2,551 17,421 Required 1. Calculate each companys net profit margin expressed as a percent. (Round your answers to 1 decimal place.) Expedia Priceline Net Profit Margin 2. Which company has generated a greater return of profit from each revenue dollar? Expedia Priceline
Answer #1
1) Expedia Priceline
Net profit margin Net income/Total revenues 10.84% 27.66%
=723/6672 =2551/9224
2) Select "Priceline"